OpenStudy (anonymous):

I am trying to solve this problem by elimination. I have one part correct but am having problem with the second part. The problem is 5x+5y=-13 7x=3y=17

OpenStudy (anonymous):

Is the second line 7x - 3y = 17? Multiply the first equation by 7/5 to get: 7x + 7y = -91/5 7x - 3y = 17 Eliminate by subtracting equation 2 from equation 1 7x + 7y = -91/5 0 + 10y = -176/5 y = -176/50 = -3.52 Put y into original equation and solve for x. 7x - 3(-3.52) = 17 x = 23/25 = .92